OpenStudy (anonymous):

Find the equation of the line tangent to y^2(y^2-4)=x^2(x^2-5) using Implicit Differentiation

myininaya (myininaya):

\[(y^2)'=2 y y'\] \[(y^2-4)'=2yy'-0=2yy'\] \[(x^2)'=2x\] \[(x^2-5)'=2x-0=2x\] recall the product rule \[(fg)'=f'g+fg'\]

OpenStudy (anonymous):

Thanks sooo very much!